40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Thu Aug 8 21:25:36 2013, Koji, Update, General, Post-vent alignment cont'd 
    Reply  Fri Aug 9 16:49:35 2013, Jenne, manasa, Update, Electronics, Post-vent alignment cont'd - RFPDs 
Message ID: 8989     Entry time: Thu Aug 8 21:25:36 2013     Reply to this: 8990
Author: Koji 
Type: Update 
Category: General 
Subject: Post-vent alignment cont'd 

- IPANG aligned on the QPD. The beam seems to be partially clipped in the chamber.

- Oplev of the IFO mirrors are aligned.

- After the oplev alignment, ITMX Yaw oplev servo started to oscillate. Reduced the gain from -50 to -20.

ELOG V3.1.3-